OpenStudy (anonymous):

y-15=3(x-12) in slope intercept form

jimthompson5910 (jim_thompson5910):

solve for y y-15=3(x-12) y-15=3x-36 y = 3x-36+15 y = 3x - 21

jimthompson5910 (jim_thompson5910):

Basically the goal is to convert the given equation into y = mx+b form, which is close to the same thing as saying "solve for y" (but you have to do a bit of simplification first)